# DNA-DetectLLM: Unveiling AI-Generated Text via a DNA-Inspired Mutation-Repair Paradigm

### v1

@@ Abstract
The rapid advancement of large language models (LLMs) has blurred the line between AI-generated and human-written text. This progress brings societal risks such as misinformation, authorship ambiguity, and intellectual property concerns, highlighting the urgent need for reliable AI-generated text detection methods. However, recent advances in generative language modeling have resulted in significant overlap between the feature distributions of human-written and AI-generated text, blurring classification boundaries and making accurate detection increasingly challenging. To address the above challenges, we propose a DNA-inspired perspective, leveraging a repair-based process to directly and interpretably capture the intrinsic differences between human-written and AI-generated text. Building on this perspective, we introduce DNA-DetectLLM, a zero-shot detection method for distinguishing AI-generated and human-written text. The method constructs an ideal AI-generated sequence for each input, iteratively repairs non-optimal tokens, and quantifies the cumulative repair effort as an interpretable detection signal. Empirical evaluations demonstrate that our method achieves state-of-the-art detection performance and exhibits strong robustness against various adversarial attacks and input lengths. Specifically, DNA-DetectLLM achieves relative improvements of 5.55% in AUROC and 2.08% in F1 score across multiple public benchmark datasets.
@@ 4.1 Experimental Setup
We compare DNA-DetectLLM with existing training-based and training-free methods. For training-based methods, we include OpenAI-D solaiman2019release , Biscope NEURIPS2024_bc808cf2 and R-Detect song2025deep . For training-free methods, we consider classic zero-shot detectors including Likelihood hashimoto-etal-2019-unifying , LogRank gehrmann-etal-2019-gltr , and Entropy ippolito-etal-2020-automatic , along with several recent SOTA approaches such as DetectGPT pmlr-v202-mitchell23a , Fast-DetectGPT bao2024fastdetectgpt , Binoculars hans2024spotting , and Lastde++ DBLP:journals/corr/abs-2410-06072 .
@@ 4.3.1 Robustness against Various Attacks
Figure 4 illustrates the AUROC curves of DNA-DetectLLM and other baselines against various attacks (see Appendix E for details). We conducted experiments on texts generated by GPT-4 Turbo, Gemini-2.0 Flash, and Claude-3.7 Sonnet, incorporating two distinct attacks: token-level edits and paraphrasing. Editing attacks involved random insertion, deletion, or substitution of tokens at rates of 1%. The paraphrasing attacks employed DIPPER to rephrase AI-generated texts. To maintain clean labels, attacks were exclusively applied to AI-generated texts.
@@ 4.4 Ablation Studies
We further evaluated the importance of repair order and different base LLMs through two types of ablation experiments. More detailed ablation studies are available in Appendix F.
@@ Appendix C Prompt Design for Main Data
Table 3 presents the general-purpose prompts we designed for different writing tasks. Using these prompts, we generated 4,800 AI-generated texts—corresponding to human-written texts—across GPT-4 Turbo, Gemini-2.0 Flash, and Claude-3.7 Sonnet, which were used in our experiments.
Appendix D Implementation Details
@@
Appendix E Robustness Experiments
@@
Appendix F Ablation Study Supplement

### v2

@@ Abstract
The rapid advancement of large language models (LLMs) has blurred the line between AI-generated and human-written text. This progress brings societal risks such as misinformation, authorship ambiguity, and intellectual property concerns, highlighting the urgent need for reliable AI-generated text detection methods. However, recent advances in generative language modeling have resulted in significant overlap between the feature distributions of human-written and AI-generated text, blurring classification boundaries and making accurate detection increasingly challenging. To address the above challenges, we propose a DNA-inspired perspective, leveraging a repair-based process to directly and interpretably capture the intrinsic differences between human-written and AI-generated text. Building on this perspective, we introduce DNA-DetectLLM, a zero-shot detection method for distinguishing AI-generated and human-written text. The method constructs an ideal AI-generated sequence for each input, iteratively repairs non-optimal tokens, and quantifies the cumulative repair effort as an interpretable detection signal. Empirical evaluations demonstrate that our method achieves state-of-the-art detection performance and exhibits strong robustness against various adversarial attacks and input lengths. Specifically, DNA-DetectLLM achieves relative improvements of 5.55% in AUROC and 2.08% in F1 score across multiple public benchmark datasets. Code and data are available at https://github.com/Xiaoweizhu57/DNA-DetectLLM.
@@ 4.1 Experimental Setup
We compare DNA-DetectLLM with existing training-based and training-free methods. For training-based methods, we include OpenAI-D solaiman2019release , Biscope NEURIPS2024_bc808cf2 and R-Detect song2025deep . For training-free methods, we consider classic zero-shot detectors including Likelihood hashimoto-etal-2019-unifying , LogRank gehrmann-etal-2019-gltr , and Entropy ippolito-etal-2020-automatic , along with several recent SOTA approaches such as DetectGPT pmlr-v202-mitchell23a , Fast-DetectGPT bao2024fastdetectgpt , Binoculars hans2024spotting , and Lastde++ DBLP:journals/corr/abs-2410-06072 . More baseline comparisons are provided in Appendix E.
@@ 4.3.1 Robustness against Various Attacks
Figure 4 illustrates the AUROC curves of DNA-DetectLLM and other baselines against various attacks (see Appendix F for details). We conducted experiments on texts generated by GPT-4 Turbo, Gemini-2.0 Flash, and Claude-3.7 Sonnet, incorporating two distinct attacks: token-level edits and paraphrasing. Editing attacks involved random insertion, deletion, or substitution of tokens at rates of 1%. The paraphrasing attacks employed DIPPER NEURIPS2023_575c4500 to rephrase AI-generated texts. To maintain clean labels, attacks were exclusively applied to AI-generated texts.
@@ 4.4 Ablation Studies
We further evaluated the importance of repair order and different base LLMs through two types of ablation experiments. More detailed ablation studies are available in Appendix G.
@@ 5 Conclusion
Acknowledgments
This work is supported by the Postdoctoral Fellowship Program of CPSF under Grant Number GZC20251076, and the National Natural Science Foundation of China (No.U2336202).
@@ Appendix C Prompt Design for Main Data
Table 3 presents the general-purpose prompts we designed for different writing tasks. Using these prompts, we generated 4,800 AI-generated texts—corresponding to human-written texts—across GPT-4 Turbo, Gemini-2.0 Flash, and Claude-3.7 Sonnet, which were used in our experiments. To ensure reproducibility, we explicitly report the generation parameters for each API call (note that Top-$k$ is not manually configurable):
  • GPT-4 Turbo: gpt-4-turbo-2024-04-09, Temperature = 1.0, Top-$p$ = 1.0.
  • Gemini 2.0 Flash: gemini-2.0-flash-001, Temperature = 1.0, Top-$p$ = 0.95.
  • Claude 3.7 Sonnet: claude-3-7-sonnet@20250219, Temperature = 1.0, Top-$p$ = 1.0.
Appendix D Main Experiments Supplement
@@
Table 4: Comparison of F1 score across public benchmark datasets.
Method M4 DetectRL Multi-LLM DetectRL Multi-Domain RealDet Avg.
OpenAI-D 68.01 70.55 67.89 70.41 69.22
Biscope 71.75 72.00 68.91 81.23 73.97
R-Detect 67.14 66.56 66.26 67.55 66.88
Entropy 75.39 70.10 55.04 60.33 65.22
Likelihood 66.82 66.62 66.60 66.87 66.73
LogRank 66.80 66.69 66.60 66.82 66.73
DetectGPT 54.53 42.41 30.79 66.59 48.58
Fast-DetectGPT 81.27 75.84 68.06 84.72 77.47
Binoculars 84.82 80.97 76.24 82.15 81.05
Lastde++ 82.74 69.17 61.72 84.59 74.56
DNA-DetectLLM 85.15 84.49 83.94 84.72 84.58
To ensure a fairer performance comparison, we select fixed thresholds for all methods based on scores computed on a separate clean dataset (e.g., DNA-DetectLLM: 0.6533, Binoculars: 0.9366, etc.). Subsequently, we recompute and report the F1 scores in Table 4 using these fixed thresholds. The clean dataset used for threshold selection consists of over 3,000 samples generated by GPT-4, Gemini, and Claude based on human-written texts sourced from XSum, WritingPrompt, and Arxiv.
Appendix E Additional Performance Comparisons
Table 5: Comparison of AUROC (%) across benchmark datasets.
Method XSum WritingPrompt Arxiv PubMedQA Avg.
Revise-Detect 39.73 65.54 95.31 66.86
GECScore 70.84 66.31 64.91 67.35
DNA-GPT 65.46 75.22 70.13 82.32 73.28
ImBD 88.07 93.06 91.06 92.59 91.20
GPTZero 99.01 98.54 94.42 88.48 95.11
DNA-DetectLLM 99.31 98.86 95.00 97.08 97.56
As shown in Table 5, we expanded our comparative experiments to include additional baselines—Revise-Detect [40], GECScore [34], DNA-GPT [38], IMBD [6] (recent but non–state-of-the-art methods), and GPTZero (a widely used commercial detector). We also incorporated results on the biomedical short-text dataset PubMedQA [19], which further demonstrate the strong and consistent performance of DNA-DetectLLM across diverse domains and detection settings.
Appendix F Robustness Experiments
@@
Table 9: F1 score (%) under paraphrasing and editing attacks
Method GPT-4 Turbo Gemini-2.0 Flash Claude-3.7 Sonnet Avg.
Insert Deletion Substitution Paraphrase Insert Deletion Substitution Paraphrase Insert Deletion Substitution Paraphrase
OpenAI-D 57.17 58.23 58.91 74.24 59.52 60.24 60.34 72.99 65.00 66.04 65.15 71.93 64.15
Biscope 54.31 55.19 50.76 27.09 86.46 86.34 86.30 42.42 86.09 85.80 85.80 32.73 64.94
R-Detect 66.46 66.46 66.46 66.50 66.16 66.16 66.16 66.16 66.63 66.63 66.63 66.63 66.42
Entropy 62.53 65.04 61.92 56.82 65.95 68.69 64.68 57.09 68.27 71.36 67.02 57.20 63.88
Likelihood 66.65 66.65 66.65 66.61 66.54 66.54 66.54 66.54 66.69 66.69 66.69 66.65 66.62
LogRank 66.61 66.61 66.61 66.61 66.57 66.57 66.57 66.50 66.65 66.65 66.65 66.65 66.60
DetectGPT 21.80 30.87 17.13 21.04 54.86 65.84 52.22 30.18 43.03 54.86 38.79 22.13 37.73
Fast-DetectGPT 77.10 84.28 76.24 89.38 87.90 89.04 87.85 88.03 77.00 85.22 75.81 84.80 83.55
Binoculars 81.33 87.16 81.15 91.50 92.44 93.08 92.28 91.65 83.15 88.71 82.36 88.78 87.80
Lastde++ 60.86 80.67 60.39 87.33 82.46 87.16 81.88 85.87 61.46 82.22 61.26 83.02 76.22
DNA-DetectLLM 86.74 90.36 86.01 93.09 94.22 94.94 93.91 93.63 87.28 91.58 86.81 91.06 90.80
@@
Table 9 reports the F1-score performances for the primary robustness experiments. Notably, DNA-DetectLLM consistently exhibits superior robustness compared to other baselines in practical scenarios.
Appendix G Ablation Study Supplement